### v2

# Disentangling Core and Edge Mechanisms of the L-mode Density Limit in DIII-D

Abstract
The L-mode density limit is re-examined on the DIII-D tokamak using a negative triangularity configuration to study the underlying physics. We find the limit may not be a singular disruptive boundary but a multifaceted density saturation phenomenon governed by distinct core and edge transport mechanisms. Sustained, non-disruptive operation is achieved at densities up to 1.8 times the Greenwald limit ($n_{\mathrm{G}}$) until the termination of auxiliary heating. Systematic power scans reveal distinct power scalings for the core ($n_{e}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) and edge ($n_{e}\propto P_{\mathrm{SOL}}^{0.42\pm 0.04}$) density limits. The edge density saturation is triggered abruptly by the onset of a non-disruptive, high-field side radiative instability that clamps the edge density below $n_{\mathrm{G}}$. In contrast, the core density continues to rise until it saturates at a level featured by substantially enhanced core turbulence. Core transport evolves from a diffusive to an intermittent, avalanche-like state, as indicated by heavy-tailed probability density functions (kurtosis $\approx 6$), elevated Hurst exponents, and a $1/f$-type power spectrum. These findings demonstrate that the L-mode density limit is determined by a combination of edge radiative instabilities and core turbulent transport. This distinction provides separate targets for control strategies aimed at extending the operational space of future fusion devices.
I Introduction
High plasma density is a fundamental requirement for achieving a self-sustaining fusion reaction, as fusion power scales strongly with density [1]. However, tokamak operation is constrained by an operational density limit [2, 3, 4]. Exceeding this critical density leads to a rapid loss of confinement, often culminating a plasma termination. Overcoming this long-standing obstacle is therefore a critical step toward realizing economically viable fusion energy.
The conventional metric for the density limit is the empirical Greenwald scaling [2, 3]: $n_{\text{G}}\,(10^{20}\,\mathrm{m^{-3}})=\frac{I_{\text{p}}\,(\text{MA})}{\pi a^{2}\,(\mathrm{m^{2}})}$, where $I_{\text{p}}$ denotes the plasma current and $a$ the minor radius. This scaling is notably independent of heating power. This paradigm is increasingly challenged by observations that the achievable density scales favorably with heating power, particularly in the low-confinement (L-mode) regime. Recent theoretical modeling [5, 6] and database analyses [7, 8] suggest that even in L-mode, the achievable density limit may scale positively with increased heating power, potentially exceeding the conventional Greenwald density. Furthermore, the power-dependent density limits have also been observed in stellarators [9, 10], which lack plasma current, suggesting a more universal physics based on power balance that transcends specific magnetic configurations. Therefore, further dedicated investigations are needed to develop a physics-based understanding of the fundamental mechanisms governing density limits in magnetically confined plasmas.
The underlying physics of the L-mode density limit involves a complex interplay of several processes. Turbulent transport [11, 12, 13, 14, 15], radiative power losses [16, 7, 4], and magnetohydrodynamic (MHD) instabilities [17, 18] have all been reported as key contributors. These processes do not act in isolation; instead, they are coupled synergistically [3] and ultimately constrain the plasma density. Consequently, developing predictive models of density limits requires disentangling their individual roles and interdependencies.
To elucidate the fundamental physics of the L-mode density limit, it is advantageous to use a plasma configuration that approximates L-mode edge conditions over a wide range of operational parameters. This avoids the complexities introduced by the L-H transition, edge transport barriers (ETBs), and edge localized modes (ELMs), which are characteristic of conventional H-mode plasmas. The negative triangularity (NT) plasmas [19, 20], characterized by an inverted “D”-shaped cross-section (see Fig. 1), provides an ideal platform for this study. These plasmas exhibit improved core confinement [21, 22, 23, 24] while maintaining L-mode-like edge conditions. While the NT edge is not identical to that of a conventional positive triangularity L-mode, its lack of a strong H-mode ETBs and ELMs makes it an excellent proxy for isolating the core and edge transport phenomena that govern density limits without confounding factors.
This paper presents a re-examination of the underlying physics governing the L-mode density limit using NT plasmas on the DIII-D tokamak. We report the achievement of sustained, non-disruptive high-density operation, where the plasma density can surpass the conventional Greenwald density by a factor of 1.8 at substantial input power. A systematic power scan reveals disparate scaling dependencies for the core ($n_{e}\propto P^{0.27\pm 0.03}_{\text{SOL}}$) and the separatrix ($n_{e}\propto P^{0.42\pm 0.04}_{\text{SOL}}$) regions. Our detailed experimental measurements reveal that different mechanisms determine density limits in the core and edge plasmas. At the plasma edge, the density is constrained substantially by enhanced turbulence driven by radiative instabilities and edge cooling. In contrast, the core density is not directly limited by the radiative cooling but eventually saturates due to the emergence of enhanced, avalanche-like turbulence. These findings challenge the traditional paradigm of a single disruptive limit, advancing a more nuanced picture where distinct processes govern density saturation in different plasma regions.
The remainder of this paper is structured as follows. Section II describes the experimental arrangement, including the plasma configuration and the diagnostic suite employed in this study. Section III presents our experimental findings, organized into four subsections: (1) high-density operation beyond the Greenwald density and the power dependence of the core and edge density limits; (2) the temporal evolution of equilibrium profiles and turbulence approaching the density limit; (3) the edge density limit and the role of radiative edge cooling; (4) the core density limit and the dominant role of avalanche-like turbulence. Finally, Section IV summarizes the key findings and discusses their broader implications for fusion research.
II Experimental Arrangement
The experiment was conducted on the DIII-D tokamak [25], with auxiliary heating power varied provided by neutral beam injection (NBI) and varied between discharges. All discharges employed a lower single-null configuration with an open divertor geometry, and the ion $\mathbf{B}\times\nabla\mathbf{B}$ drift was directed away from the primary divertor, as illustrated in Fig. 1. The plasma shape was characterized by negative triangularity at both the top and bottom ($\delta_{\text{top}}=-0.3$ and $\delta_{\text{bot}}=-0.62$), moderate elongation ($\kappa\approx 1.3$), and an edge safety factor of $q_{95}=4.1$. This paper focuses specifically on discharges with toroidal field $B_{t}=2.0$ T and plasma current $I_{p}=0.6$ MA. Additional experiments exploring different plasma current and toroidal field combinations will be reported separately [26]. In each discharge, plasma density was gradually increased via gas puffing from the top of the mainchamber. Consistent with the behavior of NT plasmas, no edge localized modes were observed in any discharge.
Figure 1: Cross-section of DIII-D tokamak showing locations of key diagnostics used for the present study.
A comprehensive suite of diagnostics systems monitored the evolving plasma state, including its equilibrium profiles, radiated power, and turbulent fluctuations. Figure 1 shows the locations of the key diagnostics. Electron density and temperature in the main chamber were measured using the Thomson scattering (TS) system [27]. Charge exchange recombination (CER) spectroscopy [28] measured the temperature, density, and toroidal and poloidal rotation velocities of carbon ions. These impurity ion profiles were used to compute radial electric field profiles ($E_{r}$) using the carbon ion force balance equation. Bolometer arrays [29] provided measurements of the radiated power distribution and total radiated power.
Several diagnostics were used to probe density fluctuations. A multichannel $\text{CO}_{2}$ interferometer [30] measured the line-averaged density and its low-$k$ fluctuations along a horizontal midplane chord and three vertical chords at major radii of $R=1.48\,\mathrm{m}$ (high-field side), $R=1.94\,\text{m}$ (core region), and $R=2.10\,\text{m}$ (low-field side). The radial interferometer polarimeter (RIP) [31] also provided line-integrated measurements of electron density fluctuations at the mid-plane. A 2D array of high-sensitivity beam emission spectroscopy (BES) [32] measured ion-scale ($k_{\theta}<3\,\mathrm{cm^{-1}}$) density fluctuations at the edge. Two 8-channel Doppler back-scattering (DBS) systems [33, 34] measured turbulent flows and low-to-intermediate-$k$ ($2<k_{\theta}<10\,\mathrm{cm^{-1}}$) density fluctuations. The radial locations and corresponding wavenumbers of the DBS measurements were calculated using the 3D ray-tracing code GENRAY [35].
III Results
III.1 Achieving High Density Beyond the Greenwald Limit
III.1.1 Density Saturation Beyond the Greenwald Limit
An important result of this study is demonstration of the sustained L-mode plasma operation at densities exceeding the conventional Greenwald density limit. This is shown in Fig. 2, which presents the time evolution of key parameters in two representative discharges with different levels of auxiliary heating from NBI. Across the experiment, the NBI power was varied from 3 MW up to 13 MW. As shown in Figs. 2(b) and 2(c), both core and separatrix densities increase with input power. The core density at $\rho\approx 0.7$, which is similar to the line-averaged density, consistently exceeds the Greenwald density ($n_{\text{G}}$). Crucially, the discharges with medium to high power NBI ($P_{\text{NBI}}>4$ MW) did not disrupt upon reaching their maximum density. Instead, they entered a phase of sustained saturated density that persisted until the termination of auxiliary heating.
Figure 2: Time history of relevant plasma parameters in two representative NT discharges with different power input levels. (a) NBI power input, (b) core density ($\rho=0.7$), (c) separatrix density, (d) the radiated power in the divertor, and (e) radiated power in the core or mainchamber. The dashed horizontal line in (b) and (c) indicates the Greenwald density.
The radiated power in the divertor (Fig. 2(d)) and the main chamber (Fig. 2(e)) also increase with density initially, indicating enhanced radiative losses. The spikes in the divertor radiation power (Fig. 2(d)) are associated with the development of a radiation front that propagates from the divertor towards the high-field side (HFS).
The line-averaged density, inferred from Thomson scattering profiles, reached $1.8\times n_{\text{G}}$ in the high-power discharge. For comparison, the CO2 interferometer measured a line-averaged density of $2.0\times n_{\text{G}}$ [20, 26]. This discrepancy arises because the interferometer’s measurement includes the substantial scrape-off layer density that is presented during high-density operation. Therefore, Thomson scattering measurements are used for the analysis in this study.
III.1.2 Distinct Power Scaling for Core and Edge Density Limit
To explore the relationship between input power and achievable density, we examine the dependence of the maximum density on the power flowing into the scrape-off layer (SOL), $P_{\mathrm{SOL}}$. Here, $P_{\mathrm{SOL}}=P_{\text{in}}-P_{\text{rad,core}}-\frac{\mathrm{d}W}{\mathrm{dt}}$, accounting for total input power, core radiated power, and changes in stored energy. Averaged values over a 100 ms window are used.
Figure 3: The maximum density achieved at two different radial positions (core and separatrix) in negative triangularity (NT) discharges as a function of the loss power entering the scrape-off layer ($P_{\mathrm{SOL}}$). The horizontal dashed line indicates the Greenwald density.
Figure 3 shows different scaling relationships for the core ($\rho=0.7$) and at the separatrix. The density at $\rho=0.7$ is chosen to represent the core region, as it is close to the line-averaged density measured by the Thomson scattering. The maximum core density increases with a relatively weak power dependence, scaling as $n_{e,\text{core}}\propto P^{0.27\pm 0.03}_{\mathrm{SOL}}$. This is consistent with L-mode density limit scalings observed on other devices such as ASDEX Upgrade [8]. In contrast, the density near the separatrix exhibits a stronger dependence, scaling as $n_{e,\text{sep}}\propto P^{0.42\pm 0.04}_{\mathrm{SOL}}$. This power dependence is reminiscent of the power-dependent density limit observed in stellarators [9, 10], suggesting a more universal physical basis involving power balance and transport.
These disparate scalings are a key finding, as they imply that distinct physical processes govern density saturation in the core and edge. The remainder of this paper will investigate these processes, proposing that turbulence dominates the core density limit, while radiative dynamics are critical at the edge.
III.2 Temporal Evolution of Profiles Approaching the Density Limit
III.2.1 Profiles Evolution Towards Density Limit
Figure 4: Temporal evolution of equilibrium profiles in a high-power discharge (shot 194133). (a) Electron density $n_{e}$, (b) electron temperature $T_{e}$, (c) carbon ion temperature $T_{i}$, and (d) toroidal rotation rate $\omega_{\phi}$. The profiles are colored according to the time in milliseconds (ms), with earlier times shown in yellow and later times shown in purple. The dashed vertical line indicates the location of the separatrix.
To investigate the underlying physical mechanisms, the temporal evolution of key plasma profiles is examined as each discharge approaches the density limit. Figure 4 displays the evolution of electron density ($n_{e}$), electron and ion temperature ($T_{e}$, $T_{i}$), and toroidal rotation rate ($\omega_{\phi}$) for a high-power discharge. As the overall density is ramped up (indicated by the color change from yellow to purple), several critical changes emerge.
While the density profile (Fig. 4(a)) rises across the plasma radius, it remains relatively flat in the core, indicating persistent particle transport. Concurrently, both the electron and ion temperature profiles (Figs. 4(b) and 4(c)) decrease substantially. A prominent feature is the flattening of the $T_{e}$ profile at the edge, a classic sign of intensified edge cooling. The toroidal rotation profile (Fig. 4(d)), which is initially peaked in the core, collapses as the density increases, leading to a significant reduction in core velocity shear.
Figure 5: The temporal evolution of the edge radial electric field ($E_{r}$) at the outer mid-plane in a high-power discharge (shot 194133). The profiles are color-coded by time, with earlier times shown in yellow and later times in purple. The vertical dashed lines indicate the approximate location of the separatrix.
Figure 5 shows the evolution of the edge radial electric field profile ($E_{r}$) at the outer mid-plane in the high-power discharge, derived from the carbon ion force balance equation. The $E_{r}$ profile initially peaks in the core region ($\rho<0.8$), but collapses progressively as the density increases. The edge $E_{r}$ profile remains relatively flat throughout the density ramp-up. This weakening of $E_{r}$ is primarily attributed to the reduced core toroidal rotation discussed previously.
These profile changes point towards a degradation in plasma confinement. The collapse of the core rotation and the associated $E_{r}\times B$ velocity shear likely weakens the turbulence suppression. This loss of flow shear facilitates the enhanced turbulent transport that ultimately limits the core density, a topic explored in subsequent sections.
III.2.2 Divergent Evolution of Turbulence in Different Regions
Figure 6: Time evolution of electron density and temperature and turbulence characteristics during a high-density discharge (shot 194133): (a) Local $n_{e}$ measured at $\rho$ = 0.7 and 0.95; (b) Local $T_{e}$ at $\rho$ = 0.7 and 0.95; (c) Spectrogram and (d) RMS amplitude of density fluctuations measured by Doppler back-scattering at the low-field side mid-plane ($\rho\approx 0.95$); (e) Spectrogram and (f) RMS amplitude of line-integrated density fluctuations measured by CO2 interferometer at the high field side; (g) Spectrogram and (h) RMS amplitude of line-integrated density fluctuations measured by RIP along the mid-plane. The vertical dashed line indicates the onset of enhanced turbulence.
As the plasma approaches density saturation, a clear divergence in turbulent behavior emerges between different plasma regions. Figure 6 illustrates this by tracking electron density and temperature alongside multichannel turbulence measurements during a high-power discharge. The vertical dashed line at approximately 3700 ms marks the onset of this divergence, which features enhanced turbulence at the HFS and the core, likely due to radiative cooling. This critical event triggers distinct turbulence responses.
Initially, the local density increases across the radius (Fig. 6(a)). Concurrently, the local intermediate-$k$ turbulence at the low-field side (LFS) mid-plane ($\rho\approx 0.95$), measured by the Doppler backscattering systems, shows a steady increase in the root-mean-square (RMS) amplitude (Figs. 6(e) and 6(f)). However, following the onset of the radiative cooling, a dramatic shift occurs: the edge density abruptly saturates (blue in Fig. 6(a)), the electron temperature drops sharply (Fig. 6(b)), and the LFS edge turbulence amplitude also saturates (Fig. 6(d)).
In contrast, line-integrated, low-$k$ density fluctuations at the HFS (Figs. 6(e) and 6(f)) and midplane (Figs. 6(g) and 6(h)) show a different evolution. The fluctuation amplitudes from both the HFS CO2 interferometer and the midplane RIP diagnostic begin to grow substantially after the onset of edge cooling. This divergent behavior—saturation of localized LFS turbulence contrasted with a continued, strong rise in line-integrated HFS and midplane fluctuations—points to a significant poloidal asymmetry in the transport. The turbulence becomes intense in the HFS region where the radiative front develops, linking the enhanced fluctuations and transport to the dynamics of the edge cooling potentially driven by radiative instabilities.
III.3 The Edge Density Limit: Role of Radiative Cooling and Instabilities
III.3.1 Development of Radiation Front at the HFS edge
The evolution of profiles and turbulence suggests the formation of a radiative front and the resulting edge cooling as a potential limiting mechanism for the edge density. To elucidate the role of this radiative instability, the analysis now focuses on the edge region to examine the dynamics of the 2D radiated power distribution.
Figure 7: Temporal evolution of the 2D distribution of radiated power during a representative high-power discharge (shot 194133), reconstructed from bolometer measurements. (a) Before density ramp-up ($t=2800$ ms), (b) enhanced divertor radiation ($t=3800$ ms), (c) radiation front leaving inner divertor ($t=4800$ ms), and (d) radiation front intensified at the HFS ($t=5800$ ms).
Figure 7 shows the temporal evolution of the radiated power distribution, reconstructed from bolometer measurements in a high-power discharge. This radiative feature is geometrically similar to a MARFE (Multifaceted Asymmetric Radiation From the Edge) [16, 36]. However, in these experiments, it manifests as a non-disruptive, quasi-stable radiation front that leads to enhanced fluctuations and density saturation. The intense but localized radiative edge cooling has also been observed in divertor detachment experiments in NT discharges [37].
Initially (Fig. 7(a), $t=2800$ ms), before the density ramp-up, radiation is concentrated near the inner divertor target, typical of an attached divertor regime. As the density is raised (Fig. 7(b), $t=3800$ ms), enhanced radiation emerges near the X-point, signaling the development of the radiative front. With a further density increase (Fig. 7(c), $t=4800$ ms), the condensation front becomes fully developed and moves upward along the high-field side separatrix, a process strongly correlated with the edge cooling shown previously in Fig. 6. Finally, the radiation front propagates further towards the HFS (inboard), and the radiated power near the separatrix intensifies (Fig. 7(d), $t=5800$ ms). It is noteworthy that in this discharge the disruption occurs shortly after a drop in the NBI power (at $t\approx 5800$ ms), which triggered an $n=1$ locked mode, and is not directly caused by the radiative edge cooling.
III.3.2 Edge Density Saturation by Radiative Instabilities
Previous observations suggest that the edge density saturates as radiative cooling intensifies. To quantify this relationship, we compare the evolution of the edge density with the amplitude of line-integrated density fluctuations. As the interferometer chord on the high-field side (HFS) passes directly through the developing radiation front, its signal serves as a reliable proxy for the intensity of the radiative instability.
Figure 8: Edge density at $\rho=0.95$ is plotted against the RMS amplitude of the density fluctuations measured by the HFS CO2 interferometer. The power law relationships for low- and high-turbulence regimes are indicated by the gray and cyan lines, respectively. The plot is in log-log scale. The data points are colored according to the energy confinement factor $H_{98y2}$. The dashed horizontal line represents the Greenwald density.
Figure 8 plots the edge density (at $\rho=0.95$) against the RMS amplitude of density fluctuations on a log-log scale. The data are collected from multiple discharges and colored by the energy confinement factor $H_{98y2}$, reveal a sharp bifurcation between two distinct operational regimes.
In the low-turbulence regime, characterized by high $H_{98y2}$ values, the edge density increases nearly linearly with HFS turbulence, following a power law of $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{1.01}$ (gray line in Fig. 8(a)). This indicates a phase where the increasing density turbulence does not yet limit the achievable edge density and plasma confinement.
This relationship breaks down abruptly as the HFS turbulence exceeds a critical threshold, marking a transition to a high-turbulence regime. Here, the fluctuation amplitude, now dominated by the radiative instability itself, continues to grow by nearly an order of magnitude. In stark contrast, the edge density saturates completely, exhibiting a nearly flat scaling of $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{0.1}$ (cyan line in Fig. 8(b)). Critically, this density saturation is accompanied by a sharp degradation in confinement, with $H_{98y2}$ dropping significantly. The overall scaling relationship is $n_{e,\text{edge}}\propto\tilde{n}_{\text{HFS}}^{0.26}$. This analysis suggests that the edge density limit is not a gradual process, but a sharp transition triggered by enhanced turbulence, likely driven by a radiative instability and the associated edge cooling. This instability clamps the edge density at a level consistently below the Greenwald limit and can be related to the loss of the edge confinement.
III.3.3 Turbulence Dynamics and the Breakdown of Zonal Flow Regulation
While the preceding analysis points to the radiative instability as the primary limit on edge density, it is crucial to understand the interplay between the turbulence and the developing radiation front. For this purpose, we use 2D measurements from the BES diagnostic, which provides a detailed look at ion-scale density fluctuations. This analysis utilizes a lower-power discharge, where the radiative cooling occurs at lower density, allowing for higher-quality BES signals.
Figure 9: Radial profiles of LFS edge turbulence characteristics during radiative cooling in a low-power NT discharge (shot 194131): (a) adiabaticity parameter, (b) normalized density fluctuation amplitude, (c) mean poloidal phase velocity, and (d) Reynolds stress $\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, (e) particle flux $\langle\tilde{n}\tilde{v}_{r}\rangle$, and (f) turbulent spreading flux $\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$. The profiles are colored by the time from lighter to darker. Particularly, the profiles corresponding to the onset of the radiative cooling (about 3700 ms) are marked by dashed curves in purple.
Figure 9 shows the evolution of key turbulence characteristics at the LFS edge as the radiative front forms (onset around 3700 ms, dashed purple curves). A fundamental change in the turbulence character is signaled by the dimensionless adiabaticity parameter [38, 13], $\alpha_{\text{adia}}=k_{\parallel}^{2}v_{te}^{2}/(\omega\nu_{ei})$. Here, $k_{\parallel}=1/qR$ is the parallel wavenumber, $v_{te}$ is the electron thermal velocity, $\nu_{ei}=\frac{\sqrt{2}e^{4}n_{i}\ln\Lambda}{12\pi^{3/2}\epsilon_{0}^{2}m_{e}^{1/2}T_{e}^{3/2}}$ is the electron-ion collision frequency, and $\omega$ is the turbulence frequency. As the edge cools, $\alpha_{\text{adia}}$ drops below unity, indicating a transition to a non-adiabatic regime, which is known to enhance turbulence and weaken zonal flow generation [13, 14].
Consistent with this transition, the normalized density fluctuation amplitude ($\tilde{n}/\bar{n}$, Fig. 9(b)) grows steadily. The mean poloidal phase velocity ($\bar{V}_{\theta}$, Fig. 9(c)) remains relatively unchanged until the onset of radiative cooling. Crucially, the ability of the turbulence to self-regulate via zonal flows collapses. This is evident in the Reynolds stress ($\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, Fig. 9(d)), the primary drive for zonal flows. The Reynolds stress flattens, indicating that the nonlinear transfer of energy from micro-turbulence to stabilizing macroscopic flows, $\bar{V}_{\theta}\partial_{r}\langle\tilde{v}_{\theta}\tilde{v}_{r}\rangle$, has been severely compromised.
The consequence of this weakly regulated turbulence is an increase in outward transport. The turbulent particle flux ($\langle\tilde{n}\tilde{v}_{r}\rangle$, Fig. 9(e)) is directed outward and grows, while the turbulent spreading flux ($\langle\tilde{n}^{2}\tilde{v}_{r}\rangle$, Fig. 9(f)) indicates that turbulence intensity is spreading inward from the separatrix. These detailed measurements indicate a potential causal sequence: edge cooling drives a transition to a non-adiabatic turbulent state where the zonal flow self-regulation mechanism fails. This leads to enhanced outward transport, providing the micro-physical link between the onset of strong radiative cooling and the ultimate saturation of the edge density.
III.4 The Core Density Limit: Role of Avalanche-like Turbulence
While the edge radiative condensation directly limits the edge density, the core density continues to increase, even surpassing the Greenwald density after the radiative front has formed. This divergence strongly suggests that a different, core-localized mechanism is at play. We now present evidence that this mechanism is a fundamental change in the nature of core turbulent transport.
III.4.1 Signatures of Avalanche-like Transport in the Core
The collapse of stabilizing $E_{r}\times B$ shear (as shown in Fig. 5) allows enhanced core turbulence. This is confirmed by examining line-integrated density fluctuations from the RIP diagnostic at the midplane. As the plasma approaches its density limit, the statistical character of this turbulence undergoes a significant change.
Figure 10: (a) Time evolution of the probability density function of the amplitudes of line-integrated density fluctuations measured by the RIP system at the midplane. (b) The time history of the kurtosis of the amplitudes of the line-integrated density fluctuations. The horizontal dashed line presents the kurtosis value of a Gaussian distribution.
Figure 10(a) shows the evolution of the probability density function (PDF) of these fluctuation amplitudes (high-pass filtered at $f>5$ kHz to avoid machine vibration noise). Early in the density ramp-up, the PDF is Gaussian (yellow), characteristic of diffusive, random-walk-like transport. However, as the density limit is approached—and particularly after the onset of the edge radiative cooling—the PDF develops pronounced “heavy tails” (purple). This indicates a transition to a state dominated by large-amplitude, intermittent transport events, a clear departure from simple diffusive behavior.
This transition is quantified by the kurtosis of the PDF (Fig. 10(b)), a measure of its “tailedness.” The kurtosis rises significantly after the onset of the radiative cooling and continues to increase, confirming the growing intermittency. Further evidence comes from the power spectrum of the fluctuations (Fig. 11(a)). After the radiative cooling onset, a distinctive “$1/f$-type” power-law scaling emerges at low frequencies. This type of spectrum is a recognized signature of self-organized criticality (SOC) and avalanche-like dynamics [39, 40, 41], where small perturbations can trigger cascading transport events of all sizes.
Figure 11: (a) Auto-power spectrum of the line-integrated midplane density fluctuations measured by RIP at different times. The dashed line indicates the power-law spectrum of $1/f^{0.8}$. (b) Time evolution of the Hurst exponent calculated from the RIP data (blue) and energy confinement factor (orange). The horizontal dashed line indicates a Hurst exponent of 0.5, corresponding to a random walk process.
Finally, the transition to an avalanche-like state is supported by the Hurst exponent ($H$), a measure of long-range temporal correlations in the data (Fig. 11(b)). The Hurst exponent rises from $H\approx 0.5$ (indicative of uncorrelated, random events) to $H\approx 0.8$ (indicative of persistent, correlated dynamics) as the density limit is approached. Consistently, the energy confinement factor $H_{98y2}$ also decreases substantially when the Hurst exponent $H$ begins to increase, indicating a loss of confinement associated with this transition. Collectively, the heavy-tailed PDFs, increased kurtosis, $1/f$ spectrum, and elevated Hurst exponent point to a consistent picture: core transport transitions from a diffusive to an intermittent, avalanche-like regime, which is ultimately responsible for limiting the core density.
III.4.2 Core Density Saturation by Avalanche-like Turbulence
The evidence for a transition to avalanche-like transport leads to a clear hypothesis: this enhanced turbulence is the primary factor limiting the core density. To test this hypothesis, we examine the direct relationship between the core density (at $\rho\approx 0.7$) and the amplitude of the midplane turbulence, as shown in Fig. 12. The data from multiple discharges, plotted on a log-log scale and colored by the energy confinement factor ($H_{98y2}$), reveal a distinct bifurcation between two transport regimes.
In the low-turbulence regime, corresponding to high confinement ($H_{98y2}>1.0$, yellow data points), the core density rises robustly with the turbulence amplitude, following a relatively strong power-law scaling of $n_{\text{e,core}}\propto\tilde{n}_{rms}^{0.49}$. This represents a phase where the core can still be effectively fueled, and the density increases despite the presence of turbulence.
Figure 12: Core density at $\rho=0.7$ is plotted against the RMS amplitude of the density fluctuations measured by the midplane RIP interferometers on a log-log scale (a) and linear scale (b). The data points are colored according to the normalized total radiation. The dotted horizontal line represents the Greenwald density. Power law relationships for low- and high-turbulence regimes are indicated by the gray and cyan lines, respectively.
However, as the avalanche-like transport intensifies and confinement degrades ($H_{98y2}<1.0$, purple data points), this relationship changes dramatically. The scaling becomes substantially weaker, flattening to $n_{\text{e,core}}\propto\tilde{n}_{\text{rms}}^{0.17}$. This transition to a weaker scaling indicates that the avalanche-like density fluctuations increase turbulent transport so effectively that they prevent significant increases in the core density. Notably, this saturation occurs at densities well above the Greenwald limit, reaching nearly twice the Greenwald density.
The saturation of the core density is more clearly visible in the linear-scale plot in Fig. 12(b). The overall response of the core density to the turbulence amplitude scales as $n_{\text{core}}\propto\tilde{n}_{\text{rms}}^{0.26}$, which also underscores the saturated behavior with increasing turbulence. The weak dependence on turbulence in this regime aligns well with the macroscopic power scaling of the core density limit ($n_{\mathrm{e,core}}\propto P_{\mathrm{SOL}}^{0.27\pm 0.03}$) found earlier. This alignment, combined with the avalanche-like dynamics, suggests that the core density limit is governed by a fundamental transition in the core turbulent transport itself, which ultimately saturates the density profile.
IV Conclusion and Discussion
This study has re-examined the L-mode density limit by leveraging the unique properties of negative triangularity plasmas on the DIII-D tokamak. Our findings suggest that the density limit, at least in this configuration, is not a singular disruptive boundary, but a multifaceted saturation phenomenon governed by distinct physical mechanisms: an edge density limit due to radiative edge cooling, and a core density limit driven by avalanche-like turbulent transport. Moreover, this study demonstrated sustained, non-disruptive operation at densities up to 1.8 times the Greenwald limit with substantial auxiliary heating.
At the plasma edge, the density limit manifests as an abrupt saturation event. This limit is triggered by the onset of enhanced turbulence driven by a radiative instability on the high-field side. While geometrically similar to a disruptive MARFE, this radiative instability acts as a quasi-stable front that clamps the edge density at a level below the Greenwald limit, without terminating the discharge.
In contrast, the core density is not directly constrained by the edge cooling and continues to rise well beyond the Greenwald limit. Its saturation is instead governed by a fundamental shift in the nature of core turbulence. As the density increases, core transport transitions from a diffusive state to one characterized by intermittent, avalanche-like dynamics. This is supported by signatures including heavy-tailed PDFs, $1/f$-type power spectra, and elevated Hurst exponents. This highly persistent, non-diffusive transport regime ultimately prevents further increases in the core density.
The dual mechanisms of regional density saturation are consistent with the disparate power scalings observed for the core ($n_{\text{e,core}}\propto P_{\mathrm{SOL}}^{0.27}$) and edge ($n_{\text{e,edge}}\propto P_{\mathrm{SOL}}^{0.42}$). The power-dependent nature of this limit and its link to transport physics show a compelling resemblance to density limits observed in stellarators. This suggests a potentially universal physics of power balance and transport that transcends specific magnetic configurations, pointing to a more fundamental basis for the density limit in magnetically confined plasmas.
Furthermore, these results suggest that high Greenwald fraction operation is feasible in future tokamaks in negative triangularity configurations, provided that the edge cooling and core turbulence dynamics are well understood and controlled. Specifically, innovative divertor solutions should be developed to manage the edge radiative cooling, while advanced core turbulence control should be explored to optimize core density and confinement, thereby extending the operational space of future reactors.
Acknowledgments
The authors gratefully acknowledge the invaluable support and contributions of the entire DIII-D team in performing this experiment. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Fusion Energy Sciences, using the DIII-D National Fusion Facility, a DOE Office of Science user facility, under Awards Nos. DE-FC02-04ER54698, DE-SC0019352, DE-SC0019004, DE-FG02-97ER54415, DE-AC52-07NA27344, and DE-SC0016154. The author (P.H.D.) acknowledges support from LLNL-led SciDAC ABOUND Project SCW1832. The author (O.S.) performed this work within the framework of the EUROfusion Consortium, via the Euratom Research and Training Programme (Grant Agreement No. 101052200– EUROfusion) and funded by the Swiss State Secretariat for Education, Research and Innovation (SERI). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union, the European Commission, or SERI. Neither the European Union nor the European Commission nor SERI can be held responsible for them.
Disclaimer: This report was prepared as an account of work sponsored by an agency of the United States Government. Neither the United States Government nor any agency thereof, nor any of their employees, makes any warranty, express or implied, or assumes any legal liability or responsibility for the accuracy, completeness, or usefulness of any information, apparatus, product, or process disclosed, or represents that its use would not infringe privately owned rights. Reference herein to any specific commercial product, process, or service by trade name, trademark, manufacturer, or otherwise does not necessarily constitute or imply its endorsement, recommendation, or favoring by the United States Government or any agency thereof. The views and opinions of authors expressed herein do not necessarily state or reflect those of the United States Government or any agency thereof.